gas occupies 900.0 mL at a temperature of 27.0 °C. What is the temperature if the same gas now has a volume of 1.75 L?


1
Expert's answer
2021-01-27T04:53:14-0500

By using Charles Law,

"\frac{V_1}{T_1}=\frac{V_2}{T_2}"

Given in question, "V_1=900\ ml,T_1=27\degree C=(273+27)K=300K"

"V_2=1.75\ L=1750\ ml, T_2=?"

Substituting these values,

"\frac{900}{300}=\frac{1750}{T_2}\implies T_2=583.33\ K"

Or, "T_2=(583.33-273)\degree C=310.33\degree C"